Fermi National Laboratory


Accelerator Update for October 21, 2005 - October 24, 2005

Friday October 21
The day shift began with the Tevatron (TeV) in shot setup, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, with NuMI off due to chiller work, and with MiniBooNE taking beam.

Operations established Recycler only store 4462 at 9:48 AM with an initial luminosity of 120.67E30. The Antiproton Source (Pbar) resumed stacking, and the Meson MCenter experiment resumed taking beam.

Operations established beam to the Meson MTest experiment at 10:12 AM.

Operations began preparing the Recycler at 11:32 AM to stash antiprotons. The stash ended and Pbar resumed stacking at 2:43 PM. (The reason this stash took so long was that Pbar experts conducted an AP1 transfer line study during the stash.)

A Booster expert began a 1.5 hour 400 MeV area study at 3:53 PM. This study caused two periods of no beam.

MCenter and MTest resumed taking beam at 4:12 PM.

The Recycler began stashing at 8:15 PM.

At 8:35 PM, Operations reported a vacuum cascade in the Main Injector (MI) P1 transfer line. Operators recovered the vacuum system. (A vacuum cascade temporarily halts beam while operators close vacuum valves, restart vacuum pumps, and then open vacuum valves.)

The stash ended and Pbar resumed stacking at 9 PM. MCenter and MTest also resumed taking beam.

At 11:39 PM, Operations reported another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system.

Saturday October 22
The midnight shift began with Operations monitoring store 4462 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MCenter, and MTest taking beam.

The Recycler began preparing at 1:12 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 2:15 AM.

At 4:43 AM, Operations reported another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system.

At 5:20 AM, store 4462 aborted due to a TeV RF station (TRF1) trip. There was no quench. Operators reset TRF1 and reset TRF2, which had tripped off earlier.

Operations prepared CDF and DO for collision hall access at 5:59 AM.

The Recycler began preparing at 6:29AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 7:28 AM.

Both CDF and D0 were out of access by 7:54 AM.

Operations put the TeV into a dry squeeze at 8:21 AM.

Operations put the TeV into shot setup at 8:53 AM.

At 9:17 AM, Operations reported yet another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system. This held off beam for about ten minutes.

Operations established Recycler only store 4464 at 10:55 AM with an initial luminosity of 138.81E30. Pbar, MCenter, and MTest resumed normal operations.

The Recycler began preparing at 1:29 PM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 2:35 PM.

The Recycler began preparing at 7:09 PM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 8:21 PM.

Sunday October 23
The midnight shift began with Operations monitoring store 4464 and the stash, with Pbar off due to the stash, and with MiniBooNE and NuMI taking beam.

The Recycler began preparing at 12:24 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 1:21 AM.

At 4:49 AM, Operations reported another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system.

The Recycler began preparing at 5:25 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 6:25 AM.

At 8:17 AM, Operations reported having problems with Booster transition. Operators investigated.

The Recycler began preparing at 10:31 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 11:26 AM.

Operations terminated store 4464 at 5:03 PM and then, ten minutes later, put the TeV into shot setup.

Operations established store 4465 at 7:22 PM with an initial luminosity of 144.512E30. This store set a New Fermi and World Luminosity Record.

At 7:23 PM, Operations reported another vacuum cascade in the MI P1 transfer line. Operators recovered the vacuum system.

The Recycler began preparing at 7:36 PM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 7:59 PM.

Monday October 24
The midnight shift began with Operations monitoring store 4465 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began preparing at 12:07 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 1:05 AM.

The Recycler began preparing at 5:15 AM to stash. The stash ended and Pbar, MTest, and MCenter resumed normal operations by 6:14 AM.

The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar and MI experts to conduct studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/24/2005   email Fermilab